OpenStudy (anonymous):

f(x)=8x√-3 Find the domain of the function

OpenStudy (anonymous):

can you rewrite the function?

OpenStudy (anonymous):

What do you mean?

OpenStudy (anonymous):

you wrote \[f(x)=8x\sqrt{-3}\] which is probably a typo of some sort

OpenStudy (anonymous):

No, that is the problem

OpenStudy (anonymous):

perhaps it was \[f(x)=8x\sqrt{x-3}\] or something similar

OpenStudy (anonymous):

\(\sqrt{-3}\) is not a real number

OpenStudy (anonymous):

YES! Sorry you are right

OpenStudy (anonymous):

and i am assuming this is not a class in complex variables

OpenStudy (anonymous):

Sorry oops

OpenStudy (anonymous):

no worries,it is \[f(x)=8x\sqrt{x-3}\]???

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

then you have to make sure that \(x-3\geq 0\) and therefore the domain is \(x\geq 3\)